# Reconstructing and resampling: a guide to utilising posterior samples from gravitational wave observationsA

### v1

@@ 4.2 Changing the waveform model for GW150914
To illustrate the application of resampling to gravitational-wave observations, we take the GWTC-2.1 data released for the LVK analysis of GW150914 Abbott et al. (2024); LIGO Scientific Collaboration and Virgo Collaboration (2022) with the IMRPhenomXPHM waveform model (Pratten et al., 2021). We then reconstruct the likelihood, using the open data from GWOSC (Abac et al., 2025b), and then calculate a new likelihood replacing the waveform with IMRPhenomXP (a special case of IMRPhenomXPHM which does not contain higher-order modes). This is in contrast to how resampling would usually be used (see, e.g Payne et al., 2019), but we provide it as a simple illustration of the techniques rather than a scientific result. In particular, the IMRPhenomXP primary pass posterior distribution is wider than the posterior found using the IMRPhenomXPHM model. This is expected, since the higher-order modes contain additional information about the source parameters that isn’t captured by the dominant mode alone.
@@
Figure 7: The posterior distribution (presented using a Gaussian KDE) on the primary mass of GW150914 as presented in the GWTC-2.1 analysis using IMRPhenomXPHM (orange dashed curve). We compare this with the posterior distributions for the IMRPhenomXP waveform calculated from resampling the GWTC-2.1 results with PSRS (pink dash-dotted curve) and PSIS (greeen dotted curve). We also include a new re-analysis using identical settings to GWTC-2.1, but with the IMRPhenomXP model (blue solid curve). For the resampled posteriors, we repeat the resampling multiple times and plot the median and 90% intervals to illustrate the variance in the result.
@@ 4.3 Changing the PSD for GW150914
As another example of the application of resampling, we again consider the GWTC-2.1 re-analysis of GW150914. For that analysis, BayesWave was used to calculate an on-source PSD. However, the BayesWave algorithm has since been upgraded (Gupta & Cornish, 2024), for example, using Akima splines rather than cubic splines. We use this upgraded version to calculate a new PSDs. In Fig.˜8, we compare the ASD (amplitude spectral density: the square root of the PSD) for the two LIGO detectors that observed the event, illustrating that while the overall shape is consistent, there are small differences between the two.
Resampling provides an easy way to check the impact these differences may have on the scientific results. We recalculate the likelihood under the new PSDs and then resample from the initial sample set in Fig.˜9, showing the detector-frame (i.e. redshifted) chirp mass. We select this parameter for visualisation as it is the best measured value. We find that the change in PSD results in a few-percent shift in the posterior, consistent with the findings of Biscoveanu et al. (2020). We also find that the resampling is efficient, with a good Pareto-smoothing diagnostic and efficiency for both PSRS and PSRS.
Finally, we calculate the Bayes factor between the new BayesWave PSD and the original analysis to be,
$\ln{\mathcal{B}}=$42.0$\pm$0.01$\$\,, (37)
suggesting the updated PSD provides a significantly better fit to the data.
@@ 5 Discussion
In this work, we have provided a comprehensive guide to reconstructing likelihoods and priors from gravitational-wave posterior samples and applying resampling techniques to generate new posterior distributions under modified assumptions. This capability is fundamental to many analyses in gravitational-wave astronomy, from population studies to tests of general relativity, yet the practical implementation details have not been systematically documented until now. Our analysis demonstrates that accurate reconstruction of the likelihood and prior is achievable when the computing environment and analysis configuration are properly matched. The residual differences we observe (Fig.˜1) are typically at the level of one part in a thousand or better for the likelihood, which we have shown is sufficiently accurate to have negligible impact on resampling efficiency. This level of precision requires careful attention to software versions, data processing details, and numerical implementation choices. The hardware-dependent variations we identified highlight an important but often overlooked aspect of reproducible gravitational-wave analysis. While exact byte-level reproduction requires identical hardware, the small random shifts introduced by different floating-point implementations are typically negligible for scientific applications, which should reassure researchers that resampling can be performed reliably across different computing environments.
@@
We also introduce Pareto-smoothing, demonstrating that it can provide moderate improvements in the efficiency and comes with an automated diagnostic which provides a useful indicator of when Pareto-smoothing can be trusted.
@@ Acknowledgements
We provide a data release associated with this paper (Ashton, 2025), which contains the Python programs to recreate the results. This includes programs to reconstruct the likelihood and priors from LVK public data releases, and programs to resample the distribution for the given examples. We include a lightweight Python module which implements the various resampling algorithms discussed within that can be easily ported or modified to users’ projects. If using this program (with modification), we ask authors to reference this work.
@@ Appendix A Probability-probability tests
The PP test is a standard diagnostic tool used to assess the reliability of Bayesian posterior distributions in gravitational-wave parameter estimation (Veitch et al., 2015; Romero-Shaw et al., 2020). The test compares the cumulative distribution of confidence intervals from a set of simulated injections against the uniform distribution expected for unbiased posteriors. We follow the Bilby implementation of the PP test and provide a simplified version in the data release notebooks. Specifically, we take the result from injection studies and compute the fraction of injected parameter values that fall below various quantile levels of their corresponding posterior distributions. For each parameter of interest, the PP test calculates the cumulative probability $P(\theta<\theta_{\rm sim}$) where $\theta_{\rm sim}$ is the simulated value and the probability is computed using the recovered posterior samples. If the posteriors are unbiased and properly calibrated, these $p$-values should be uniformly distributed between 0 and 1. The PP plot displays the empirical cumulative distribution function of these $p$-values against the theoretical uniform distribution, with deviations from the diagonal indicating systematic biases in the parameter estimation. We include a Kolmogorov-Smirnov test to compare the empirical distribution of $p$-values against the uniform distribution, providing a $p$-value that indicates the probability of observing such deviations by chance. In the plot, we also include confidence bands around the diagonal using the beta distribution to account for finite sample size effects. The width of these bands depends on the number of injections, with larger injection campaigns providing tighter constraints on systematic biases.
@@
In the main text, Section 4, we use the PP test to demonstrate that the resampling methods can unbias a posterior given a corrected posterior density. The specific case given considers over- and under-constrained posteriors generated by varying the ratio of $\sigma_{s}$ (the standard deviation of the simulated measurement error) to $\sigma_{p}$ (the standard deviation of the simulated posteriors). Specifically, in Fig.˜5, we consider the case when the posteriors are too narrow $\sigma_{s}/\sigma_{p}=2$. This results in a characteristic ‘S’ shape. This is one possible characteristic failure mode of the PP plot. In Fig.˜10, we extend this to study the other three failure modes: a case where the posterior is under-constrained (producing an inversion of the ‘S’); biased positively (producing a bow-like deviation above the diagonal) and and negatively (producing a bow-like deviation below the diagonal). We include these figures to help users diagnose bias from resampling (or from direct sampling) approaches.
@@
We note that this behaviour is specific to PP tests which use quantile levels to calculate the fraction of injected parameters that fall outside the posterior at a given level. If instead, the highest-posterior density is used, the behaviour will differ.

### v2

@@ 4.2 Changing the waveform model for GW150914
To illustrate the application of resampling to gravitational-wave observations, we take the GWTC-2.1 data released for the LVK analysis of GW150914 Abbott and others (2024); LIGO Scientific Collaboration and Virgo Collaboration (2022) with the IMRPhenomXPHM waveform model (Pratten and others, 2021). We then reconstruct the likelihood, using the open data from GWOSC (Abac and others, 2025f), and then calculate a new likelihood replacing the waveform with IMRPhenomXP (a special case of IMRPhenomXPHM which does not contain higher-order modes). This is in contrast to how resampling would usually be used (see, e.g Payne et al., 2019), but we provide it as a simple illustration of the techniques rather than a scientific result. In particular, the IMRPhenomXP primary pass posterior distribution is wider than the posterior found using the IMRPhenomXPHM model. This is expected, since the higher-order modes contain additional information about the source parameters that isn’t captured by the dominant mode alone.
@@
Figure 7: The posterior distribution (presented using a Gaussian KDE) on the primary mass of GW150914 as presented in the GWTC-2.1 analysis using IMRPhenomXPHM (orange dashed curve). We compare this with the posterior distributions for the IMRPhenomXP waveform calculated from resampling the GWTC-2.1 results with PSRS (pink dash-dotted curve) and PSIS (green dotted curve). We also include a new re-analysis using identical settings to GWTC-2.1, but with the IMRPhenomXP model (blue solid curve). For the resampled posteriors, we repeat the resampling multiple times and plot the median and 90% intervals to illustrate the variance in the result.
@@ 4.3 Changing the PSD for GW150914
As another example of the application of resampling, we again consider the GWTC-2.1 re-analysis of GW150914. For that analysis, BayesWave (Littenberg and Cornish, 2015) was used to calculate a PSD, using data surrounding the signal and modelling the noise properties with a parameterised model including cubic splines and Lorentzian lines. However, the version of BayesWave used to produce LVK parameter estimates has since been upgraded (Gupta and Cornish, 2024), for example, using Akima splines rather than cubic splines. We use this upgraded version to calculate a new PSD with settings similar to those used in the GWTC-4.0 analysis (Abac and others, 2025c). In Fig.˜8, we compare the ASD (amplitude spectral density: the square root of the PSD) for the two LIGO detectors that observed the event, illustrating that while the overall shape is consistent, there are small differences between the two.
Resampling can provide an easy way to check the impact these differences may have on the scientific results. However, we have to be careful in our choice of likelihood since the PSD enters Eq.˜3 in a normalisation term that is often ignored (cf. Thrane and Talbot, 2019). This correction is not included in the Bilby likelihood computation by default, and so we add it explicitly (see the data release for details). We calculate the likelihood and likelihood ratio under the old and new PSD then using the difference in log-likelihoods, resample from the initial sample set. In Fig.˜9, we plot the distribution of the detector-frame (i.e. redshifted) chirp mass. We select this parameter for visualisation as it is the best measured value. We find that the change in PSD results in a few-per-cent shift in the posterior, consistent with the findings of Biscoveanu et al. (2020). We also find that the resampling is efficient, with a good Pareto-smoothing diagnostic and efficiency for both PSRS and PSRS.
Now turning to the Bayes factor, from Eq.˜5, we see that three different Bayes factors can be computed when changing from the “old” (GWTC-2.1) to our “new” PSD (using the GWTC-4.0 settings). First, the ratio of the signal evidence, second, the ratio of the noise evidence, and third, the ratio of the signal vs noise evidence. We tabulate these in Table˜3. The first row, the Bayes factor of signal evidence, indicates a strong preference for the old PSD over the new PSD. However, this comparison includes information about both how well the signal model explains the data and how well the PSD explains the noise properties of the data. We can separate these by considering the second row, the ratio of noise evidence, which indicates an even stronger preference for the old PSD. Finally, when we compute the ratio of the signal vs noise Bayes factors, we effectively ask “Which PSD makes the signal stand out more strongly?” and now we find evidence in favour of the new PSD. This suggests the new PSD, while assigning a lower overall probability to the data, provides a noise model against which the signal is more distinguishable. We do not investigate this point further, but note that Bayes factors are not typically used as a means to determine the choice of PSD: for details of the development of the BayesWave see Gupta and Cornish (2024). Moreover, we highlight that the differences could be related to the settings we chose for the PSD generation.
Bayes factor Estimate
$\ln\left(\frac{\mathcal{Z}(d|M,P_{\rm new})}{\mathcal{Z}(d|M,P_{\rm old})}\right)$ $1684.86$\pm$0.02$$
$\ln\left(\frac{\mathcal{Z}(d|N,P_{\rm new})}{\mathcal{Z}(d|N,P_{\rm old})}\right)$ 1726.90-1726.90
$\ln\left(\frac{\mathcal{Z}(d|M,P_{\rm new})}{\mathcal{Z}(d|N,P_{\rm new})}\frac{\mathcal{Z}(d|N,P_{\rm old})}{\mathcal{Z}(d|M,P_{\rm old})}\right)$ $42.04$\pm{0.02}$
Table 3: Estimates of the three log Bayes factors that can be computed from the study changing the PSD in the analysis of GW150914. For the first and last lines, the evidences are computed from Eq.˜29 using different choices of the likelihood or likelihood ratio. Meanwhile, for the second row, the ratio of noise evidence, there is no uncertainty, as this is computed directly from the data and PSD. Note that all three are computed directly from the likelihoods themselves, but the final row can also be computed as the difference in the first two rows.
@@ 5 Discussion
In this work, we have provided a comprehensive guide to reconstructing likelihoods and priors from gravitational-wave posterior samples and applying resampling techniques to generate new posterior distributions under modified assumptions. This capability is fundamental to many analyses in gravitational-wave astronomy, from population studies to tests of general relativity, yet the practical implementation details have not been systematically documented until now. Our analysis demonstrates that accurate reconstruction of the likelihood and prior is achievable when the computing environment and analysis configuration are properly matched. The residual differences we observe (Fig.˜1) are typically at the level of one part in a thousand or better for the likelihood, which we have shown is sufficiently accurate to have a negligible impact on resampling efficiency. This level of precision requires careful attention to software versions, data processing details, and numerical implementation choices. The hardware-dependent variations we identified highlight an important but often overlooked aspect of reproducible gravitational-wave analysis. While exact byte-level reproduction requires identical hardware, the small random shifts introduced by different floating-point implementations are typically negligible for scientific applications, which should reassure researchers that resampling can be performed reliably across different computing environments.
@@
We also introduce Pareto-smoothing, demonstrating that it can provide moderate improvements in efficiency and comes with an automated diagnostic which provides a useful indicator of when Pareto-smoothing can be trusted.
@@ Acknowledgements
We provide a data release associated with this paper (Ashton, 2026), which contains the Python programs to recreate the results. This includes programs to reconstruct the likelihood and priors from LVK public data releases, and programs to resample the distribution for the given examples. We include a lightweight Python module which implements the various resampling algorithms discussed within that can be easily ported or modified to users’ projects. If using this program (with modification), we ask authors to reference this work.
@@ Appendix A Probability-probability tests
The PP test is a standard diagnostic tool used to assess the reliability of Bayesian posterior distributions in gravitational-wave parameter estimation (Veitch and others, 2015; Romero-Shaw and others, 2020). The test compares the cumulative distribution of confidence intervals from a set of simulated injections against the uniform distribution expected for unbiased posteriors. We follow the Bilby implementation of the PP test and provide a simplified version in the data release notebooks. Specifically, we take the result from injection studies and compute the fraction of injected parameter values that fall below various quantile levels of their corresponding posterior distributions. For each parameter of interest, the PP test calculates the cumulative probability $P(\theta<\theta_{\rm sim}$) where $\theta_{\rm sim}$ is the simulated value and the probability is computed using the recovered posterior samples. If the posteriors are unbiased and properly calibrated, these $p$-values should be uniformly distributed between 0 and 1. The PP plot displays the empirical cumulative distribution function of these $p$-values against the theoretical uniform distribution, with deviations from the diagonal indicating systematic biases in the parameter estimation. We include a Kolmogorov-Smirnov test to compare the empirical distribution of $p$-values against the uniform distribution, providing a $p$-value that indicates the probability of observing such deviations by chance. In Fig.˜10, we also include confidence bands around the diagonal using the beta distribution to account for finite sample size effects. The width of these bands depends on the number of injections, with larger injection campaigns providing tighter constraints on systematic biases.
@@
In the main text, Section 4, we use the PP test to demonstrate that the resampling methods can unbias a posterior given a corrected posterior density. The specific case given considers over- and under-constrained posteriors generated by varying the ratio of $\sigma_{s}$ (the standard deviation of the simulated measurement error) to $\sigma_{p}$ (the standard deviation of the simulated posteriors). Specifically, in Fig.˜5, we consider the case when the posteriors are too narrow $\sigma_{s}/\sigma_{p}=2$. This results in a characteristic ‘S’ shape. This is one possible characteristic failure mode of the PP plot. In Fig.˜10, we extend this to study the other three failure modes: a case where the posterior is under-constrained (producing an inversion of the ‘S’); biased positively (producing a bow-like deviation above the diagonal), and negatively (producing a bow-like deviation below the diagonal). We include these figures to help users diagnose bias from resampling (or from direct sampling) approaches.
@@
We note that this behaviour is specific to PP tests, which use quantile levels to calculate the fraction of injected parameters that fall outside the posterior at a given level. If instead, the highest-posterior density is used, the behaviour will differ.